6. Swathmore Clothing Corporation grants its customers 30 days’ credit. The company uses the allowance method...

6.

Swathmore Clothing Corporation grants its customers 30 days’ credit. The company uses the allowance method for its uncollectible accounts receivable. During the year, a monthly bad debt accrual is made by multiplying 2% times the amount of credit sales for the month. At the fiscal year-end of December 31, an aging of accounts receivable schedule is prepared and the allowance for uncollectible accounts is adjusted accordingly.

At the end of 2020, accounts receivable were $604,000 and the allowance account had a credit balance of $68,000. Accounts receivable activity for 2021 was as follows:

Beginning balance $ 604,000
Credit sales 2,770,000
Collections (2,633,000 )
Write-offs (54,000 )
Ending balance $ 687,000

The company’s controller prepared the following aging summary of year-end accounts receivable:

Summary
Age Group Amount Percent Uncollectible
0−60 days $ 445,000 4 %
61−90 days 72,000 12
91−120 days 64,000 23
Over 120 days 106,000 38
Total $ 687,000

Required:
1. Prepare a summary journal entry to record the monthly bad debt accrual and the write-offs during the year.
2. Prepare the necessary year-end adjusting entry for bad debt expense.
3-a. What is total bad debt expense for 2021?
3-b. How would accounts receivable appear in the 2021 balance sheet?
Answer #1

SOLUTION

1.

S.No. Accounts titles and Explanation Debit ($) Credit ($)
1. Bad debt expense 55,400
Allowance for uncollectible accounts 55,400
(To record monthly bad debts expense)
2. Allowance for uncollectible accounts 54,000
Accounts receivable 54,000
(To record write offs of receivable)

Bad debt expense = Credit sales * 2%

= $2,770,000 * 2% = 55,400

2.

S.No. Accounts titles and Explanation Debit ($) Credit ($)
1. Bad debt expense 12,040
Allowance for uncollectible accounts 12,040
(To record adjusting entry for bad debt expense)

Working-

Age Group Amount ($) (A) Percent Uncollectible (B) Amount Uncollectible (A*B)
0-60 days 445,000 4% 17,800
61-90 days 72,000 12% 8,640
91-120 days 64,000 23% 14,720
Over 120 days 106,000 38% 40,280
Total 687,000 81,440

Ending balance before adjustment = Beginning balance of accounts receivable + Bad debt expense - Write off

= 68,000 + 55,400 - 54,000 = 69,400

Bad debt expense at year end = Allowance for doubtful accounts - Ending balance before adjustment

= 81,440 - 69,400 = 12,040

3a. Total bad debts = Bad debts during the year + Bad debt expense at year end

= 55,400 + 12,040 = 67,440

3b.

Assets Amount ($)
Current assets:
Accounts receivable (net) (687,000 - 81,440) 605,560